# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

### v2

@@ Abstract
The linear speedup property is essential for demonstrating the advantage of distributed algorithms over their single-node counterparts. In this paper, we study the stochastic Push-Pull method, a widely adopted decentralized optimization algorithm over directed graphs (digraphs). Unlike methods that rely solely on row-stochastic or column-stochastic mixing matrices, Push-Pull avoids nonlinear correction and has shown superior empirical performance across a variety of settings. However, its theoretical analysis remains challenging, and the linear speedup property has not been generally established—revealing a significant gap between empirical success and limited theoretical understanding. To bridge this gap, we propose a novel analysis framework and prove that Push-Pull achieves linear speedup over arbitrary strongly connected digraphs. Our results provide the comprehensive theoretical understanding for stochastic Push-Pull, aligning its theory with empirical performance. Code: https://github.com/pkumelon/PushPull.
@@ 1 Introduction
Scaling machine learning tasks to large datasets and models necessitates efficient distributed computation across multiple nodes. In distributed optimization, a network of $n$ nodes cooperates to solve
$\displaystyle\min_{x\in\mathbb{R}^{d}}\quad f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)\quad\mbox{where}\quad f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F(x;\xi_{i})].$ (1)
Here, $\xi_{i}$ denotes a random data sample drawn from the local distribution $\mathcal{D}_{i}$, and $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is the local loss function, which may be nonconvex. During optimization, each node $i$ can access its own local gradient $\nabla F_{i}(x;\xi_{i})$, but must communicate to obtain gradient information from the other nodes. Data heterogeneity is typically present in (1), meaning the local distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ differ across nodes. Consequently, the loss functions $f_{i}(x)$ and $f_{j}(x)$ are generally not equal, i.e., $f_{i}(x)\neq f_{j}(x)$.
Parallel stochastic gradient descent (SGD) is a widely adopted approach in distributed optimization, where local gradients are globally averaged using either a central server [1] or Ring All-Reduce communication primitives [2]. However, such global synchronization incurs substantial bandwidth overhead and high latency as the number of nodes increases, thereby hurting the scalability of the optimization process. To address these limitations, decentralized SGD has recently been proposed to solve problem (1) without requiring global gradient aggregation. In this setting, each node independently maintains and updates its model by communicating only with immediate neighbors, significantly reducing communication overhead. Beyond efficiency, decentralized algorithms offer improved robustness to node and link failures, maintaining convergence as long as the underlying communication graph remains connected. In contrast, centralized SGD breaks down if the central server crashes.
In this paper, we consider a general decentralized setting where the network topology is modeled as a strongly connected directed graph (digraph), which naturally encompasses undirected networks as a special case. Compared to undirected graphs, digraphs are particularly well-suited for ad hoc decentralized systems, such as heterogeneous robotic networks that operate under asynchronous, one-sided communication and random node activation. More recently, digraphs have shown great potential in distributed deep learning, where certain classes of directed networks exhibit both topological sparsity and strong algebraic connectivity. These properties make them especially attractive for the distributed training of large-scale models. Notable examples include the exponential graph [3], EquiTopo [4], and B-ary tree [5].
@@ 1.1 Decentralized algorithms over digraphs
In any connected network, the topology can be characterized by a mixing matrix. This matrix reflects how nodes are connected and is important for understanding how the network affects algorithm performance. For undirected networks, it is straightforward to construct symmetric and doubly stochastic mixing matrices. For directed networks, achieving a doubly stochastic matrix is generally infeasible. Consequently, mixing matrices in such cases are typically either column-stochastic [6, 7] or row-stochastic [8, 9], but not both.
Building on the connectivity pattern of the underlying network topology, decentralized optimization algorithms over directed graphs can be broadly categorized into three families in the literature. The first family relies on the Push-Sum technique [10, 11], which employs column-stochastic mixing matrices for all communications. Accordingly, we refer to this class as Push-Only algorithms in this paper. When exact gradients $\nabla f_{i}$ are available, the well-known Subgradient-Push algorithm [12, 11] converges to the desired solution, albeit with relatively slow sublinear rates even under strong convexity. More advanced methods such as EXTRA-Push [13], DEXTRA [14], ADD-OPT [15], and Push-DIGing [7] improve convergence speed by mitigating the effects of data heterogeneity. Stochastic variants, including SGD-Push [16] and S-ADDOPT [17], extend these approaches to settings with stochastic gradients.
The second family consists of algorithms that use only row-stochastic matrices for communication, referred to as the Pull-Only setting. Just as Push-Sum underpins the design of Push-Only algorithms, the Pull-Diag gossip protocol [9] serves as the foundation for Row-Only methods. Building on Pull-Diag, reference [9] adapted the distributed gradient descent algorithm to the Row-Only setting. Subsequently, [18, 19] extended gradient tracking techniques to this framework, while momentum-based Pull-Only gradient tracking methods were developed by [20, 21].
The third family comprises algorithms that alternate between row-stochastic and column-stochastic mixing matrices, commonly referred to as the Push-Pull method [22] or the AB method [23]. In this framework, a row-stochastic matrix is used to exchange model parameters, while a column-stochastic matrix is employed to share gradient information. As demonstrated in [23, 22, 24], the Push-Pull family typically achieves significantly faster convergence rates than algorithms relying solely on column- or row-stochastic matrices. [25] extended the Push-Pull method to stochastic optimization settings, while [5] improved its convergence rate on B-ary tree network topologies. Accelerated or constrained variant of the Push-Pull method have also been explored in [26, 27].
@@ 1.2 Linear speedup in iteration complexity
Linear speedup refers to the property of a distributed algorithm whereby the number of iterations required to reach an $\epsilon$-accurate solution decreases linearly with the number of nodes. For example, centralized SGD achieves an $\epsilon$-accurate solution in $\mathcal{O}\left(\textstyle\frac{1}{n\epsilon^{2}}\right)$ iterations , where $n$ denotes the number of nodes. This scaling behavior demonstrates that centralized SGD achieves linear speedup with respect to the number of nodes. The linear speedup property is crucial for establishing the advantage of distributed algorithms over their single-node counterparts. Without this property, the practical utility of a distributed algorithm may be limited, as it would offer little benefit from using multiple computing nodes.
The linear speedup property of decentralized algorithms over undirected graphs is well established. The first such analysis for decentralized SGD was presented in [28], demonstrating that decentralized SGD can achieve the same asymptotic convergence rate as its centralized counterpart. Building on this foundation, a variety of advanced decentralized algorithms—such as Exact Diffusion/NIDS [29, 30] and Gradient Tracking methods [31, 32, 33]—are proven to achieve linear speedup [34, 29, 35]. In addition, [36] provides a linear speedup analysis for decentralized SGD under time-varying undirected network topologies.
Recent literature has increasingly focused on establishing the linear speedup property of decentralized algorithms over digraphs. For Push-Only algorithms that employ column-stochastic mixing matrices, linear speedup has been formally proven in [16, 37]. Moreover, a recent study [38] proposes methods that achieve the optimal iteration complexity in this setting. For Pull-Only algorithms that employ row-stochastic mixing matrices, convergence has been extensively studied under various settings [9, 18, 19, 20, 21]; however, these analyses do not establish the linear speedup property. The main challenge lies in the fact that using a row-stochastic matrix alone introduces a deviation in the descent direction from the globally averaged gradient. With a novel analysis framework addressing this issue, the linear speedup property of Pull-Only algorithms was established only recently in [39]. These results reveal the effectiveness of Push-Only and Pull-Only algorithms over digraphs.
However, although Push-Pull algorithms typically achieve significantly better empirical performance than Push-Only and Pull-Only methods, their linear speedup property has not been generally established. Existing theoretical analyses of Push-Pull remain limited, primarily focusing on strongly convex loss functions in deterministic settings [23, 40, 22, 5]. In the nonconvex and stochastic setting, neither linear speedup nor even basic convergence guarantees have been formally established for Push-Pull. The most relevant theoretical result to date is presented in [5], which introduces a variant called B-ary Tree Push-Pull (BTPP). While BTPP is shown to achieve linear speedup, the result relies on the specialized B-ary tree topology, and the analysis does not generalize to general network topologies.
@@ 1.3 Motivation and main results
Building on the above discussion, the existing theoretical convergence analyses for Push-Pull methods remain rather limited, underscoring a clear discrepancy between the weak theoretical guarantees and their strong empirical performance. This gap gives rise to the following two open questions:
@@
    Can Push-Pull achieve linear speedup over a broader class of network topologies—for instance, all strongly connected digraphs?
Addressing these two questions is crucial for establishing a theoretical foundation that supports the effectiveness of Push-Pull methods.
Regarding Question Q1, we demonstrate that both Push-Only and Pull-Only algorithms closely align with centralized SGD; specifically, after a careful numerical transformation, their variable updates closely resemble those of centralized SGD. This structural similarity explains why the linear speedup property can be readily established for Push-Only and Pull-Only methods. In contrast, Push-Pull is not designed to align with centralized SGD and consistently exhibits a non-vanishing deviation from the centralized trajectory. Consequently, the linear speedup analysis for Push-Pull is substantially more challenging and fundamentally distinct from that of Push-Only and Pull-Only algorithms.
For Question Q2, we establish a linear speedup convergence result for the Push-Pull method. To the best of our knowledge, this is the first such result for Push-Pull over all strongly connected digraphs under standard assumptions, whereas existing analyses of stochastic Push-Pull [5, 41] are restricted to tree-structured topologies. By introducing a novel analysis framework that addresses the misalignment between Push-Pull and centralized SGD, this paper bridges the gap between the convergence theory and the strong empirical performance of the Push-Pull method.
@@
The remainder of the paper is organized as follows. Section 2 introduces necessary notations, assumptions, and measures. Section 3 reviews the Push-Only, Pull-Only, and Push-Pull methods, and offers insights into the derivation of linear speedup convergence. In Section 4, we present the key components of our analysis and state the main convergence theorem. Section 5 provides extensive numerical experiments that demonstrate the ability of the Push-Pull method to achieve linear speedup. Finally, Section 6 concludes the paper.
2 Notations, Assumptions and Measures
@@
This section introduces notations, assumptions, and measures that will be used throughout the paper.
@@
Notations. The directed network is denoted by $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}=\{1,2,\ldots,n\}$ is the set of nodes and $\mathcal{E}$ is the set of directed edges. An edge $(j,i)\in\mathcal{E}$ indicates that node $j$ can send information to node $i$. A directed graph is said to be strongly connected if there exists a directed path from any node to every other node in the graph. We say that $\mathcal{G}=(\mathcal{N},\mathcal{E})$ is the underlying topology of the weight matrix $A=[a_{ij}]$ if $A$ is entrywise nonnegative and $a_{ij}>0$ if and only if $(j,i)\in\mathcal{E}$. Let $\mathds{1}\in\mathbb{R}^{n}$ denote the all-ones vector. Set $[n]$ denotes $\{1,2,\ldots,n\}$. For a vector $v$, $\mathrm{Diag}(v)$ denotes the diagonal matrix with $v$ as its diagonal entries.
@@
We use matrix $A$ to denote the row-stochastic matrix (i.e., $A\mathds{1}=\mathds{1}$) and matrix $B$ to denote the column-stochastic matrix (i.e., $B^{\top}\mathds{1}=\mathds{1}$). The Perron vectors [42] of matrices $A$ and $B$, if they exist, are denoted by $\pi_{A}$ and $\pi_{B}$, respectively. We introduce $A_{\infty}:=\mathds{1}\pi_{A}^{\top}$ and $B_{\infty}:=\pi_{B}\mathds{1}^{\top}$. Additionally, we define $n\times d$ matrices
$\displaystyle{\mathbf{x}}^{(k)}$ $\displaystyle:=[({x}_{1}^{(k)})^{\top};({x}_{2}^{(k)})^{\top};\cdots;({x}_{n}^{(k)})^{\top}]$
$\displaystyle{\mathbf{g}}^{(k)}$ $\displaystyle:=[\nabla F_{1}({x}_{1}^{(k)};\xi_{1}^{(k)})^{\top};\cdots;\nabla F_{n}({x}_{n}^{(k)};\xi_{n}^{(k)})^{\top}]$
$\displaystyle\nabla f({\mathbf{x}}^{(k)})$ $\displaystyle:=[\nabla f_{1}({x}^{(k)}_{1})^{\top};\cdots;\nabla f_{n}({x}^{(k)}_{n})^{\top}]$
[…]
@@
Graphs ${\mathcal{G}}_{A}$ and ${\mathcal{G}}_{B}$ are strongly connected.
@@
Besides, there exists constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ such that
$\|A^{k}-A_{\infty}\|_{2}\leq p_{A}\lambda_{A}^{k},\quad\|B^{k}-B_{\infty}\|_{2}\leq p_{B}\lambda_{B}^{k},\quad\forall k\geq 0.$
@@
Finally, we state standard assumptions regarding the stochastic gradients and function smoothness.
@@
$\displaystyle\mathbb{E}_{k}[\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(k)})\quad\mbox{and}\quad\mathbb{E}_{k}[\|\nabla F_{i}(x_{i}^{(k)};\xi_{i}^{(k)})-\nabla f_{i}(x_{i}^{(k)})\|_{F}^{2}]\leq\sigma^{2}$
for any $i\in[n]$ and $k\geq 0$.
@@
Measures. Various effective metrics have been proposed to capture the impact of digraphs on decentralized stochastic optimization. Let $\Pi_{A}:=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}:=\mathrm{Diag}(\pi_{B})$. Given a column-stochastic mixing matrix $B$, references [25, 38] introduce $\beta_{B}:=\|\Pi_{B}^{1/2}(B-B_{\infty})\Pi_{B}^{-1/2}\|_{2}\in[0,1)$ as a measure of graph connectivity, and define $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ to reflect the deviation of the equilibrium distribution $\pi_{B}$ from the uniform distribution $n^{-1}\mathds{1}$. Similarly, reference [38] introduces $\beta_{A}:=\|\Pi_{A}^{1/2}(A-A_{\infty})\Pi_{A}^{-1/2}\|_{2}\in[0,1)$ and $\kappa_{A}:=\max\pi_{A}/\min\pi_{A}$ to characterize the influence of the row-stochastic mixing matrix $A$ on decentralized algorithms.
However, this paper introduces a new set of metrics to facilitate the convergence analysis of the Push-Pull method. When Proposition 1 holds, we define the following constants:
$\displaystyle s_{A}$ $\displaystyle:=\max\{\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}^{2}\},$ (2)
$\displaystyle s_{B}$ $\displaystyle:=\max\{\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2}^{2}\},$ (3)

### v3

@@ Abstract
The linear speedup property is essential for demonstrating the advantage of distributed algorithms over their single-node counterparts. This paper studies the stochastic Push-Pull method, a widely adopted decentralized optimization algorithm for directed graphs (digraphs). Unlike methods that rely on Push-Only or Pull-Only communications, Push-Pull avoids nonlinear correction and has shown superior empirical performance across various settings. Despite its success, whether Push-Pull achieves linear speedup has remained an open question, revealing a significant gap between its empirical performance and theoretical understanding. To bridge this gap, we propose a novel multi-step descent analysis framework and prove that Push-Pull indeed achieves linear speedup over arbitrary strongly connected digraphs. Our results provide a comprehensive theoretical understanding of the stochastic Push-Pull method, finally aligning its theoretical convergence with empirical performance. Code: https://github.com/pkumelon/PushPull.
@@ 1 Introduction
Scaling modern machine learning tasks to massive datasets and increasingly large models requires efficient distributed computation across multiple nodes. In distributed optimization, a network of $n$ nodes cooperatively solves
$\displaystyle\min_{x\in\mathbb{R}^{d}}\quad f(x):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)\quad\mbox{where}\quad f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F_{i}(x;\xi_{i})].$ (1)
where each local function $f_{i}$ is defined by a data distribution $\mathcal{D}_{i}$. Each node has access only to stochastic gradients $\nabla F_{i}(x;\xi_{i})$ from its local data, while communication is required to aggregate information across the network. In practice, the distributions $\{\mathcal{D}_{i}\}_{i=1}^{n}$ are heterogeneous, so the local losses $f_{i}$ generally differ. This heterogeneity poses significant challenges to both the design and analysis of distributed algorithms.
A natural baseline for solving (1) is centralized stochastic gradient descent (SGD), where gradients are aggregated globally either through a central server [1] or collective communication primitives such as Ring All-Reduce [2]. While centralized SGD enjoys strong theoretical guarantees—including the celebrated linear speedup property in iteration complexity—its reliance on global averaging leads to significant communication overhead and creates a single point of failure at the server. These drawbacks limit scalability in large, heterogeneous systems.
Decentralized SGD has emerged as a compelling alternative [3, 4]. In this setting, each node maintains its own local model and communicates only with its neighbors. This localized communication substantially reduces bandwidth usage and alleviates latency bottlenecks. Moreover, decentralized methods are inherently robust to failures: as long as the communication graph remains connected, the network can continue to train without dependence on a single master node.
In this paper, we consider decentralized optimization over directed graphs (digraphs). Directed communication naturally models many real-world systems, including heterogeneous robotic swarms with asymmetric links, sensor networks with one-sided message passing, and distributed deep learning environments where communication is constrained by bandwidth asymmetry. Directed networks are also attractive because certain classes of sparse digraphs, such as exponential graphs [5], EquiTopo [6], and B-ary trees [7], achieve both scalability and strong connectivity.
@@ 1.1 Decentralized algorithms over digraphs
In a connected network, the topology can be characterized by a mixing matrix. This matrix reflects how nodes are connected and is important for understanding how the network affects algorithm performance. For undirected networks, it is straightforward to construct symmetric and doubly stochastic mixing matrices. For directed networks, achieving a doubly stochastic matrix is generally infeasible. Consequently, mixing matrices in such cases are typically either column-stochastic [8, 9] or row-stochastic [10, 11], but not both.
Building on the connectivity pattern of the underlying network topology, decentralized optimization algorithms over directed graphs can be broadly categorized into three families in the literature. The first family relies on the Push-Sum technique [12, 13], which employs column-stochastic mixing matrices for all communications. Accordingly, we refer to this class as Push-Only algorithms in this paper. When exact gradients $\nabla f_{i}$ are available, the well-known Subgradient-Push algorithm [8, 13] converges to the desired solution, albeit with relatively slow sublinear rates even under strong convexity. More advanced methods such as EXTRA-Push [14], DEXTRA [15], ADD-OPT [16], and Push-DIGing [9] improve convergence rate by mitigating the effects of data heterogeneity. Stochastic variants, including SGD-Push [5] and S-ADDOPT [17], extend these approaches to settings with stochastic gradients.
The second family consists of algorithms that use only row-stochastic matrices for communication, referred to as the Pull-Only methods. Just as Push-Sum underpins the design of Push-Only algorithms, the Pull-Diag gossip protocol [11] serves as the foundation for Pull-Only methods. Reference [11] adapted the distributed gradient descent algorithm to the Pull-Only setting. Subsequently, gradient tracking techniques were extended to this framework by [18, 19], while momentum-based Pull-Only gradient tracking methods were developed in [20, 21].
The third family comprises algorithms that alternate between row-stochastic and column-stochastic mixing matrices, commonly referred to as the Push-Pull method [22] or the AB method [23]. In this framework, a row-stochastic matrix is used to exchange model parameters, while a column-stochastic matrix is employed to share gradient information. As demonstrated in [23, 22, 24], the Push-Pull family typically achieves significantly faster convergence rates than algorithms relying solely on column- or row-stochastic matrices. Reference [25, 26] extended the Push-Pull method to stochastic optimization settings, while [7] improved its convergence rate on B-ary tree network topologies. Accelerated or constrained variant of the Push-Pull method have also been explored in [27, 28].
@@ 1.2 Linear speedup in iteration complexity
The linear speedup is a central property for distributed optimization. It states that the number of iterations required to reach an $\epsilon$-accurate solution decreases linearly with the number of nodes $n$. Centralized SGD achieves this scaling, converging in $\mathcal{O}(1/(n\epsilon^{2}))$ iterations. Without this property, the benefit of parallelization is limited, as increasing the number of nodes would not yield proportional gains in iteration complexity.
For decentralized algorithms on undirected networks, linear speedup is well established. The first rigorous analysis for decentralized SGD was given in [29]. Building on this work, a range of advanced algorithms—including Exact Diffusion/NIDS [30, 31] and gradient tracking methods [32, 33, 34, 9]—have been proven to achieve linear speedup [35]. Time-varying network extensions are also addressed in [36].
For digraphs, theoretical understandings are less explored. For Push-Only algorithms with column-stochastic mixing, linear speedup has been established in [5, 37], with subsequent refinements achieving optimal iteration complexity in [38]. For Pull-Only algorithms with row-stochastic mixing, convergence has been extensively analyzed under various settings [11, 18, 19, 20, 21], but these studies did not establish linear speedup because row-stochastic matrices inherently introduce a deviation from the globally averaged gradient. Only recently, linear speedup for Pull-Only algorithms was proven in [39]. These advances reveal the effectiveness of both Push-Only and Pull-Only methods over digraphs.
Despite this progress, establishing the linear speedup property for Push-Pull algorithms—those combining both row- and column-stochastic matrices—has remained an open problem. While Push-Pull methods consistently outperform Push-Only and Pull-Only variants in practice, their theoretical understanding remains limited. Existing analyses focus mostly on strongly convex settings [23, 25, 22, 7], and provide no guarantees for nonconvex or stochastic optimization. The most relevant result to date is [7], which demonstrates linear speedup for Push-Pull under a specialized B-ary tree topology, but the analysis cannot be extended to general network structures. A recent work [40] establishes guarantees for Push-Pull in a specific non-convex federated learning setting, but does not prove the linear speedup property. Table 1 provides a summary of existing results on decentralized algorithms over digraphs.
@@ 1.3 Motivation and main results
The above discussion highlights a striking gap between theory and practice: Push-Pull algorithms consistently outperform Push-Only and Pull-Only methods in empirical studies, yet their scalability theory is far less developed. This gap motivates two fundamental questions:
@@
    Can push–pull achieve linear speedup on general strongly connected digraphs under standard assumptions?
In this work, we address both questions. Regarding Question Q1, we show that Push-Only and Pull-Only can be reformulated to align closely with centralized SGD, making their linear speedup analysis relatively straightforward. In contrast, Push-Pull exhibits a persistent, non-vanishing deviation from centralized SGD dynamics, rendering the conventional single-step descent analysis framework ineffective.
For Question Q2, we introduce a novel multi-step descent analysis framework that aggregates errors over blocks of iterations. This framework uncovers a telescoping structure in the gradient-tracking recursion: intermediate gradient noise cancels out, while boundary terms are exponentially attenuated by mixing dynamics. Consequently, the residual error remains bounded at a constant scale, enabling us to establish the desired $\mathcal{O}(1/(n\epsilon^{2}))$ iteration complexity for Push-Pull.
To the best of our knowledge, this paper presents the first general proof that Push-Pull achieves linear speedup on arbitrary strongly connected digraphs in the stochastic nonconvex setting. In particular, our multi-step descent analysis framework not only bridges the gap between theory and practice for Push-Pull but also introduces techniques that may be of independent interest for the broader study of decentralized algorithms. A summary of existing results is provided in Table 1.
Table 1: Convergence rate comparison of representative decentralized algorithms in non-convex and stochastic settings over directed graphs. “Conv. Rate” refers to the asymptotic rate as $T\to\infty$; “Iter. Complex.” denotes the number of iteration requried to achieve $\epsilon$-accuracy; “S.C.” denotes “Strongly Connected”; “R-Sto.”, “C-Sto.” and “R/C Alt.” refer to “Row-Stochastic”, “Column-Stochastic” and “Row/Column Alternate”, respectively. Note that References [25, 26] study the Push-Pull algorithm under strongly convex objectives, rather than the non-convex setting. $\sigma$ denotes gradient noise.
Algorithms
Conv.
Rate
Iter.
Complex.
Topology
Mixing
Matrix
Centralized SGD $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ Star graph -
Pull-Diag-GT [39] $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ S.C. digraph R-Sto.
Push-Diging [38] $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ S.C. digraph C-Sto.
Push-Pull [25] - $\;\;\;$-\;\; S.C. digraph R/C Alt.
Push-Pull [26] - $\;\;\;$-\;\; S.C. digraph R/C Alt.
Push-Pull [40] $\mathcal{O}(\frac{\sigma}{\sqrt{T}})$ $\mathcal{O}(\frac{\sigma^{2}}{\epsilon^{2}})$ S.C. digraph R/C Alt.
Push-Pull [7] $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ B-ary Tree R/C Alt.
Push-Pull (Ours) $\mathcal{O}(\frac{\sigma}{\sqrt{nT}})$ $\mathcal{O}(\frac{\sigma^{2}}{n\epsilon^{2}})$ S.C. digraph R/C Alt.
@@
The remainder of the paper is organized as follows. Section 2 introduces the notations, assumptions, and performance metrics used throughout the paper. Section 3 reviews the Push-Only, Pull-Only, and Push-Pull methods and provides insights into the derivation of their linear speedup properties. Section 4 presents our novel multi-step descent analysis framework and explains the intuition behind its role in enabling linear speedup analysis. Section 5 formally develops the supporting lemmas and states the main convergence theorem. Section 6 reports extensive numerical experiments that validate the ability of the Push-Pull method to achieve linear speedup. Finally, Section 7 concludes the paper.
2 Notations, Algorithms, Assumptions and Metrics
@@
This section introduces notations, assumptions, and metrics that will be used throughout the paper.
@@
Notations. The directed network is denoted by $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where $\mathcal{N}=\{1,2,\ldots,n\}$ is the set of nodes and $\mathcal{E}$ the set of directed edges. An edge $(j,i)\in\mathcal{E}$ indicates that node $j$ can send information to node $i$. The graph is strongly connected if a directed path exists between every pair of nodes. We say $\mathcal{G}$ is the underlying topology of a nonnegative weight matrix $A=[a_{ij}]$ if $a_{ij}>0$ precisely when $(j,i)\in\mathcal{E}$. Let $\mathds{1}\in\mathbb{R}^{n}$ denote the all-ones vector and $[n]:=\{1,2,\ldots,n\}$. For a vector $v$, $\mathrm{Diag}(v)$ denotes the diagonal matrix with $v$ on its diagonal, while $\min v$ and $\max v$ denote its smallest and largest entries, respectively.
@@
We let $A$ denote a row-stochastic matrix ($A\mathds{1}=\mathds{1}$) and $B$ a column-stochastic matrix ($B^{\top}\mathds{1}=\mathds{1}$), with Perron vectors [41] $\pi_{A}$ and $\pi_{B}$, respectively. Define $A_{\infty}:=\mathds{1}\pi_{A}^{\top}$ and $B_{\infty}:=\pi_{B}\mathds{1}^{\top}$. Additionally, we define $n\times d$ matrices
$\displaystyle{\mathbf{x}}^{(t)}$ $\displaystyle\hskip-1.42262pt:=\hskip-1.42262pt[({x}_{1}^{(t)})^{\top}\hskip-1.42262pt;({x}_{2}^{(t)})^{\top}\hskip-1.42262pt;\cdots;({x}_{n}^{(t)})^{\top}\hskip-1.42262pt]$
$\displaystyle\hskip-1.42262pt{\mathbf{g}}^{(t)}$ $\displaystyle\hskip-1.42262pt:=\hskip-1.42262pt[\nabla F_{1}({x}_{1}^{(t)};\hskip-1.42262pt\xi_{1}^{(t)})^{\top}\hskip-1.42262pt;\hskip-1.42262pt\cdots\hskip-1.42262pt;\hskip-1.42262pt\nabla F_{n}({x}_{n}^{(t)};\xi_{n}^{(t)})^{\top}\hskip-1.42262pt]$
$\displaystyle\nabla f({\mathbf{x}}^{(t)})$ $\displaystyle\hskip-1.42262pt:=\hskip-1.42262pt[\nabla f_{1}({x}^{(t)}_{1})^{\top};\cdots;\nabla f_{n}({x}^{(t)}_{n})^{\top}$]
where subscripts index nodes and superscripts index iterations. Let ${\mathcal{F}}_{t}$ be the filtration containing all information up to ${\mathbf{x}}^{(t)}$, excluding the stochastic gradients at iteration $t$. We use $\mathbb{E}_{t}[\cdot]=\mathbb{E}[\cdot\mid{\mathcal{F}}_{t}]$ for conditional expectation. Norms are denoted as follows: $\|\cdot\|$ for the vector $\ell_{2}$-norm, $\|\cdot\|_{F}$ for the Frobenius norm, and $\|\cdot\|_{2}:=\max_{\|v\|=1}\|Av\|$ for the matrix spectral norm.
Push-Pull algorithm. The Push-Pull algorithm, initially proposed for deterministic decentralized optimization [22, 23], can be naturally extended to the stochastic gradient setting [25]:
$\displaystyle{\mathbf{x}}^{(t+1)}$ $\displaystyle=A{\mathbf{x}}^{(t)}-\alpha{\mathbf{y}}^{(t)},$ (2a)
$\displaystyle{\mathbf{y}}^{(t+1)}$ $\displaystyle=B{\mathbf{y}}^{(t)}+{\mathbf{g}}^{(t+1)}-{\mathbf{g}}^{(t)},$ (2b)
where $A$, $B$ are row-stochastic and column-stochastic matrices determined by the underlying digraph. For initialization, we let $x_{1}^{(0)}=x_{2}^{(0)}=\cdots=x_{n}^{(0)}=x^{(0)}$ and ${\mathbf{y}}^{(0)}={\mathbf{g}}^{(0)}$. The row-stochastic mixing matrix $A$ in (2a) facilitates consensus among the local variables maintained by each node, whereas the column-stochastic mixing matrix $B$ in (2b) enables the computation of a globally averaged gradient. In the absence of gradient noise, i.e., when ${\mathbf{g}}^{(t)}=\nabla f({\mathbf{x}}^{(t)})$, it has been shown in [22, 23] that ${\mathbf{x}}^{(t+1)}$ converges to ${\mathbf{x}}^{\star}=[(x^{\star})^{\top};\cdots;(x^{\star})^{\top}]\in\mathbb{R}^{n\times d}$ for strongly convex problems, where $x^{\star}$ is the solution to problem (1).
@@
Assume ${\mathcal{G}}_{A}$ and ${\mathcal{G}}_{B}$ to be strongly connected digraphs (see the definition in notations) with $n$ nodes.
@@
Besides, there exist constants $p_{A},p_{B}>0$, $\lambda_{A},\lambda_{B}\in[0,1)$ such that
$\|A^{t}-A_{\infty}\|_{2}\leq p_{A}\lambda_{A}^{t},\quad\|B^{t}-B_{\infty}\|_{2}\leq p_{B}\lambda_{B}^{t},\quad\forall k\geq 0.$ (3)
@@
We next state assumptions regarding stochastic gradients and function smoothness.
@@
$\displaystyle\mathbb{E}_{t}[\nabla F_{i}(x_{i}^{(t)};\xi_{i}^{(t)})]$ $\displaystyle=\nabla f_{i}(x_{i}^{(t)})\quad\mbox{and}\quad\mathbb{E}_{t}[\|\nabla F_{i}(x_{i}^{(t)};\xi_{i}^{(t)})-\nabla f_{i}(x_{i}^{(t)})\|_{F}^{2}]\leq\sigma^{2}$
for any $i\in[n]$ and $t\geq 0$. In addition, we assume that stochastic gradients at different nodes are independent for any $t\geq 0$ and $1\leq i<j\leq n$.
@@
Metrics. Various effective metrics have been proposed in literautre to capture the impact of digraphs on decentralized stochastic optimization. Let $\Pi_{A}:=\mathrm{Diag}(\pi_{A})$ and $\Pi_{B}:=\mathrm{Diag}(\pi_{B})$ where $\pi_{A}$ and $\pi_{B}$ are defined in Proposition 1. Given a column-stochastic mixing matrix $B$, references [25, 38] introduce $\beta_{B}:=\|\Pi_{B}^{-1/2}(B-B_{\infty})\Pi_{B}^{1/2}\|_{2}\in[0,1)$ as a measure of graph connectivity, and define $\kappa_{B}:=\max\pi_{B}/\min\pi_{B}$ to reflect the deviation of the equilibrium distribution $\pi_{B}$ from the uniform distribution $n^{-1}\mathds{1}$. Similarly, reference [38] introduces $\beta_{A}:=\|\Pi_{A}^{1/2}(A-A_{\infty})\Pi_{A}^{-1/2}\|_{2}\in[0,1)$ and $\kappa_{A}:=\max\pi_{A}/\min\pi_{A}$ to characterize the influence of the row-stochastic mixing matrix $A$ on decentralized algorithms.
This paper introduces a new set of metrics to facilitate the convergence analysis of the Push-Pull method. When Proposition 1 holds, we define:
$\displaystyle s_{A}$ $\displaystyle:=\max\big\{\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2},\sum_{i=0}^{\infty}\|A^{i}-A_{\infty}\|_{2}^{2}\big\},$ (4)
$\displaystyle s_{B}$ $\displaystyle:=\max\big\{\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2},\sum_{i=0}^{\infty}\|B^{i}-B_{\infty}\|_{2}^{2}\big\},\vskip 5.69054pt$ (5)